A radioactive Isotope Gallium – 67, used in diagnosing malignant tumor, has a biological half – life of 46.5 hours. If we start with 100mg of the isotope, how many milligrams will be left after 24 hours
1
Expert's answer
2021-10-29T00:47:02-0400

Use the half-life decay model as follows:

"A = A _{0}(\frac{1}{2})\\^\frac{-t}{h} \\= A _{0}2^\frac{-t}{h}"

Using "A _{0} = 100" and "h = 46.5", we obtain:

"100(2^\frac{-t}{46.5})"

After 24 hours:

"100(2\\^\frac{-24}{46.5})\\= 69.9245 \ milligrams"